# Personhood credentials: Artificial intelligence and the value of

### v1

@@ 2.2   Credential limits and the goals of a personhood credentialing ecosystem
Per-person rate limits. Without per-person limits on some digital activities, harmful actors can pose as multiple people to gain outsized influence on peer-to-peer platforms and in public comment processes. [^51] For example, the public comment process for the FCC’s 2017 decision on repealing net neutrality protections was compromised by nearly 18 million bot-generated comments [191]. In addition, harmful actors can evade a service’s rules—by, for example, creating many new accounts even when caught violating a service’s policies. [^52] Repeat abuse is a large tax on digital services and their legitimate users. Gao et al. estimate the market size of the underground registration ecosystem at about 4.8–128.1 million USD per year (as of 2022) [106].
@@ 3.1   Reduce the impact of sockpuppeting
Personhood credentials are a mechanism for reducing the viability and impact of sockpuppets. Regardless of whether sockpuppets are AI-powered, services that adopt PHCs can reduce malicious actors’ ability to falsely present themselves as multiple individuals on their platform by, for example, imposing account verification limits on each personhood credential. [^56] The account verification limit does not strictly need to be “one verified account per PHC” in order for PHCs to curb malicious activity: A website could decide, for instance, that it wishes for each PHC-holder to have at most three verified accounts, which could be unlinked to one another. The important property here is the ability to enforce some finite limit by counting how many times the PHC has been used for a purpose, without more specifically identifying the holder or their particular PHC. Once a user is verified to be a person, there are many ways a site could incorporate PHCs to facilitate trustworthy discourse, e.g., by boosting or labeling such accounts, or by offering users ways to screen out accounts that are not verified. [^57] Stronger distinguishability could facilitate a more trusting online community. On today’s Internet, sometimes false accusations are made that a given account is a bot [16]. Knowing another user to in fact be a person could inspire greater trust and ultimately lead to more productive conversations. For instance, some research indicates that using a profile picture with a face in it leads to a greater empathetic response from others in online discussions [157]. PHCs may help achieve a similar effect. Scenario 1 outlines one hypothetical use. Such methods are not meant to entirely rid a service of AI-generated content or AI-powered accounts, many of which are benign. [^58] Some services have established processes for the appropriate disclosure of automated accounts, such as accounts operated with AI or through other computerized systems that do not require human intervention [275]. But PHCs do offer a range of options for curbing the impact—particularly at scale—when AI’s use is deceptive. [^59] While the introduction of PHCs does not entirely solve the problem of AI-powered deception, it dramatically reduces the scale. For example, someone might still hand over their single PHC-verified dating site account to an agentic AI they direct to carry out a catfishing scam. Yet the catfisher is less likely to be successful with a single profile as opposed to many potentially targeted profiles used in parallel. Or, someone may fill their single PHC-verified social media account with misleading AI-generated content to spread disinformation. But the spreader of disinformation will not be as successful in making their misleading content go viral without amplification from other independent-seeming accounts that they also control. In both cases, if the bad actor is found to be in violation of the service’s rules once, they can be prevented from returning—a dynamic we discuss further in the following subsection.
@@
In recent years, sockpuppets have caused harm across a number of settings. On social media, sockpuppets attempt to manipulate public opinion on a topic [119, 166, 19, 161]. [^60] For a broader investigation of sockpuppets in online communities, see [152].  [^61] Some digital services describe their site rules’ rationales in terms of reducing the impact of sockpuppets. X’s Community Notes, for instance, states a requirement for accounts to have a unique phone number to “help prevent the creation of large numbers of fake or sock puppet contributor accounts that could be used for inauthentic rating” [274]. In online dating, bot-powered profiles are used for catfishing, luring users into costly and sometimes dangerous situations [62]. In formal government comment processes, bots overwhelm genuine civic input. [^62] For example, millions of fake comments were submitted to the Federal Communications Commission’s 2017 public request for comment on net neutrality. Researchers have conducted analysis on the methods used to generate such inauthentic comments, which were far more rudimentary than is possible with today’s AI [268]. Sockpuppets continue to overwhelm US government public comment processes, and few agencies have taken action to systematically filter out or catch the malicious actors perpetrating these abuses [64, 202]. Recently, a bill aiming to curb foreign and AI-powered abuse in the public comment process for the Bureau of Land Management (the American Voices in Federal Lands Act) was introduced in the US Senate [190].
@@
PHCs might also help to advance a wide range of beneficial experiments in digital governance in the public, private, and nonprofit sectors, though these are speculative and also come with risks. [^63] Scholars have argued that new democratic innovations may be especially important in the age of highly capable AI [6, 25]. For instance, AI tools might help to digitally gather and analyze public opinions from large segments of the population [98]—PHCs can help ensure that people do not use multiple personas to wield more influence and thus increase the perceived legitimacy of such processes. [^64] PHCs can help digital governance processes be more robust to claims that AI bots manipulated a particular outcome. For more on the ability to cast doubt on outcomes by invoking the specter of AI manipulation, see [225]. Moreover, PHCs can help reduce particular communities’ abilities to intentionally undermine these processes. [^65] One challenge in soliciting broad input online is that particular communities might flood a poll to try to undermine the poll’s intent. For instance, a poll to name a new Mountain Dew flavor produced leading options like “Hitler did nothing wrong” [124], forcing the organizers to take down the naming process. For more, see [219]. Looking ahead, collective input on the design and deployment of AI systems is an emerging special case for soliciting representative opinions, free from sockpuppet manipulation. [^66] Some AI providers have commissioned research on how to best channel popular input in their tools, including via “Alignment Assemblies” to better align models with democratic will [282, 86, 63, 13].
@@ 3.2   Mitigate bot attacks
Malicious actors use botnets—coordinated groups of entities meant to appear as distinct users—to carry out more effective attacks online. Sometimes, overwhelming a Web service with a number of distinct-appearing users is the point of the attack, as in distributed denial-of-service (DDoS) attacks [237]. Other times, the purpose of relying upon many distinct-appearing users is to create many independent chances of success, like circumventing a service’s rules and creating new accounts for abuse even after having been caught. As discussed in Section 1, without new mitigations, malicious actors powered by AI systems are more likely to succeed at these attacks.
@@
Personhood credentials can help reduce the impact of bot attacks by enabling service providers to suspend users whose accounts have engaged in abuse. Today, adversaries’ tactics like changing IP addresses and spoofing different browser configurations often successfully evade the restrictions in place, at least for a period [210]. Some existing software does guard against these attacks, but the developers of cybersecurity software have been locked in a perpetual cat-and-mouse game with attackers. And now, AI changes the dynamic, powering the attackers by reducing the manual work needed to get a new botnet running again after a previous one was caught. By using personhood credentials, service providers can restrict each individual to a limited number of accounts. This means that if someone breaks the rules and their account is suspended, they—identified through their personhood credential used at registration—cannot create new accounts without limit. [^67] When establishing restrictions on user activity, websites must balance the level of abuse they are willing to tolerate against the potential loss of legitimate user engagement that such limits might cause [167].
@@
Bot attacks have wide-ranging harms. Genuine users of peer-to-peer digital services are taxed when bots are not kept in check—they face increased friction and are targets of scams. [^68] For instance, the website may have particular processes—phone number verification, rate limits—that introduce friction for benign users but are primarily meant to limit abuse by repeat bad actors. Digital communities that aim to give some asset (e.g., compute credits, products for beta testing, scientific surveys) to any individual person, but only once, are often undermined by bot attacks. PHCs can serve as a valuable tool for checking who has already received the benefit, without revealing or tracking any further information. [^69] A prosaic example of such fraud is free trials for digital services, which are often serially abused [77]. This rate-limiting can also be applied to mitigate other forms of abuse that are exacerbated by rapid, AI-driven actions, such as ticket scalping, which disadvantages ordinary buyers [250, 187]. AI-powered attempts at fraudulently acquiring benefits from governments or aid-giving organizations—even when successfully defended against—can drain resources and make it more difficult to serve the targets of the aid; screening by applicants verified with PHCs may help. [^70] One risk is that screening via PHCs may reduce fraud but also make it more difficult for legitimate recipients to receive aid: Administrative overhead to apply for benefits tends to result in many people not receiving benefits that they are properly eligible for [120, 226].
@@
Looking ahead, we expect that many forms of AI-powered deceptive behavior that are especially harmful at scale can be indirectly mitigated by personhood credentials. In the vignette in Scenario 2, for example, we discussed one particular form of AI-powered deception that is harmful in isolated incidents but especially harmful at scale—impersonation. [^71] By impersonation, we mean taking on the likeness of someone in particular, as opposed to taking on a sockpuppet persona—a plausible human that could exist but does not. We dwell on this case because it nicely delineates how PHCs can and cannot directly help when it comes to AI-powered deception online: PHCs by default cannot validate a holder’s legal identity, and so a digital likeness service cannot directly check a PHC to confirm that the holder is taking on their own likeness rather than another person’s. [^72] Malicious actors may have incentives to impersonate a real person to carry out social engineering attacks or to set up a profit-making scheme (e.g., if the target of impersonation is a celebrity). However, such activities can be prohibited by digital services’ policies (and sometimes involve breaking a law [^73] Courts have sometimes ruled in favor of a “right to publicity” and “right to privacy.” Using a person’s likeness without their permission may violate these rights, depending on the circumstances [269, 101].), and if a user is ever found to be in violation of the service’s policies, PHCs can make the policies more enforceable, by suspending the PHC-linked account and disallowing future signups for that PHC. We expect that—anticipating these consequences—fewer attackers will attempt such unauthorized impersonations.
@@ 3.3   Allow verified delegation to AI agents
Scenarios 3 Accountable delegation on peer-to-peer platforms.
@@
AI systems are increasingly capable of acting autonomously [141, 264, 95, 277]. While enabling many beneficial use cases, [^74] For a description of some economically useful properties of AI agents, see [231]. Already, some humans are deferring to AI-powered solutions for navigating dating apps on their behalf [230]. this autonomy (“agenticness”) also facilitates a new form of deception: bad actors can deploy AI systems that, instead of pretending to be a person, accurately present as AI agents but pretend to act on behalf of a user who does not exist. This strategy exploits the current lack of norms around disclosure of agentic AIs, including a lack of norms around disclosing the identities of the people controlling them (often called their “principals”). [^75] In some jurisdictions there is already regulation governing bot disclosure [143]. In 2023, legislation was introduced to codify bot-disclosure federally, though it has not become law [189, 258]. Note that these bills outline requirements around disclosure of AI use, i.e., requiring that people disclose when a bot was employed to take an action or when a piece of content was generated by an AI model. The outlined disclosure requirements stop there—they do not include disclosures related to the identity of the person or organization operating the AI.
@@
Personhood credentials could offer a way to verify that AI agents are acting as delegates of real people, signaling credible supervision without revealing the principal’s legal identity. This feature could be useful in a range of settings where AI assistants have the potential to ease the frictions of using an anonymous service; Scenario 3 outlines one such use case. Should a principal fail to address harms caused by their PHC-verified AI, they risk suspension from a service. Suspension implies that they lose their ability to verify delegates for some time period, reducing their capacity to perpetrate future harms.
@@
Note how, in this case, PHCs create a form of accountability for AI agents without demanding sensitive information from principals. [^76] A large and growing body of legal scholarship explores when and how humans should be accountable for the harms caused by AI agents [87, 54, 55, 52, 56, 156, 74, 144]. Many—though by no means all [238, 100]—of these proposals involve holding humans liable for some harms caused most directly by those agents. It is beyond the scope of this article to recommend any of these approaches over others or explore the finer points of how such theories could work. We note, however, that these theories rely on the ability of someone harmed by an AI agent to sue the principal of that agent, which in turn depends on the principal being identifiable—which a PHC alone does not achieve. Even without directly identifying the principal, however, PHCs can still shift the benefits of AI agent delegation to be more positive. This suspension mechanism may effectively signal which AI agents have trustworthy principals, even when verifying AI agents through principal-linked PHCs is voluntary. Agents that remain unverified might be perceived as having reasons for not undergoing verification. [^77] This mechanism aligns with the “unraveling result” of voluntary disclosure models in economic theory [172, 173]. In scenarios where parties hold verifiable private information—such as whether an AI agent operates under a trustworthy principal—even if revealing this information is optional, agents associated with trustworthy principals have a strong incentive to disclose it. Consequently, a lack of disclosure becomes informative: in equilibrium, agents that do not disclose are effectively signaling that they do not have trustworthy principals. Some malicious activities involving autonomous AI agents may rely on hiding the fact that multiple agents are controlled by the same individual. [^78] One example of how a person might enlist unverified agents to manipulate a market: A potential buyer on an auction site might create multiple personas to submit lowball offers, in hopes of baiting the seller into an artificially low assessment of their item’s value. In a similar example, a potential seller might enlist synthetic personas to create the appearance of higher demand for their item, inducing a legitimate buyer to pay a higher price and to act with more urgency. In each case, the ability to recognize these entities as linked to the same identity—even without knowing any further details of the identity—would help to head off these risks. For an estimation of the prevalence of these practices, known as “shill-bidding,” in online markets, see [51]. For a discussion of market designs that are robust to these practices, see [145]. PHCs can help address this issue by creating a framework that links multiple AI agents to a single principal without revealing the principal’s specific identity. By doing so, PHCs make it more difficult for bad actors to conceal their network of AI agents, thereby reducing the potential for abuse that stems from undisclosed common principals.
@@
Ultimately, a fuller framework for verifying AI agents [43, 44] and their principals will likely be necessary. Sometimes, a website—or a third-party user interacting with the agent—may need to verify the AI’s specific principal, not merely that it is backed by some principal.

### v2

@@ 2.2   Credential limits and the goals of a personhood credentialing ecosystem
Per-person rate limits. Without per-person limits on some digital activities, harmful actors can pose as multiple people to gain outsized influence on peer-to-peer platforms and in public comment processes. [^51] For example, the public comment process for the FCC’s 2017 decision on repealing net neutrality protections was compromised by nearly 18 million bot-generated comments [191]. In addition, harmful actors can evade a service’s rules—by, for example, creating many new accounts even when caught violating a service’s policies. [^52] The market size of the underground ecosystem for account registration, by one estimate, is roughly 4.8–128.1 million USD per year (as of 2022) [106]. Illicit account creation and the repeat abuse it enables are taxes on digital services and their legitimate users.
@@ 3.1   Reduce the impact of sockpuppeting
Personhood credentials are a mechanism for reducing the viability and impact of sockpuppets. Regardless of whether sockpuppets are AI-powered, services that adopt PHCs can reduce malicious actors’ ability to falsely present themselves as multiple individuals on their platform by, for example, imposing account verification limits on each personhood credential. [^56] An account verification limit does not strictly need to be “one verified account per PHC” in order for PHCs to curb malicious activity: A website could decide, for instance, that it wishes for each holder to have at most three verified accounts, which could be unlinked to one another. The important property here is the ability to enforce some finite limit by counting how many times a PHC has been used for a purpose, without more specifically identifying the holder or their particular PHC. Once a user is verified to be a person, there are many ways a site could incorporate PHCs to facilitate trustworthy discourse, e.g., by boosting or labeling such accounts, or by offering users ways to screen out accounts that are not verified. [^57] Stronger distinguishability could facilitate a more trusting online community. On today’s Internet, sometimes false accusations are made that a given account is a bot [16]. Knowing another user to in fact be a person could inspire greater trust and ultimately lead to more productive conversations. For instance, some research indicates that using a profile picture with a face in it leads to a greater empathetic response from others in online discussions [157]. PHCs may help achieve a similar effect. Scenario 1 outlines one hypothetical use. Such methods are not meant to entirely rid a service of AI-generated content or AI-powered accounts, which can often be benign. [^58] Some services have established processes for the appropriate disclosure of automated accounts, such as accounts operated with AI or through other computerized systems that do not require human intervention [275]. But PHCs do offer a range of options for curbing the impact when AI’s use is deceptive—particularly at scale. [^59] While the introduction of PHCs does not entirely solve the problem of AI-powered deception, it dramatically reduces the scale. For example, someone might still hand over their single PHC-verified dating site account to an agentic AI, which they direct to carry out a catfishing scam. Yet the catfisher is less likely to be successful with a single profile as opposed to many potentially targeted profiles used in parallel. Or, someone may post misleading AI-generated content from their single PHC-verified social media account to spread disinformation. But the spreader of disinformation will not be as successful in making their misleading content go viral without amplification from other independent-seeming accounts that they also control. In both cases, if the bad actor is found to be in violation of the service’s rules, they can be prevented from returning—a dynamic we discuss further in the following subsection.
@@
In recent years, sockpuppets have caused harm across a number of settings. On social media, sockpuppets attempt to manipulate public opinion on a topic [119, 166, 19, 161]. [^60] For a broader investigation of sockpuppets in online communities, see [152].  [^61] Some digital services describe the rationale for their sites’ rules in terms of reducing the impact of sockpuppets. X’s Community Notes, for instance, states a requirement for accounts to have a unique phone number to “help prevent the creation of large numbers of fake or sock puppet contributor accounts that could be used for inauthentic rating” [274]. In online dating, bot-powered profiles are used for catfishing, luring users into costly and sometimes dangerous situations [62]. In formal government comment processes, bots overwhelm genuine civic input. [^62] For example, millions of fake comments were submitted to the Federal Communications Commission’s 2017 public request for comment on net neutrality. Researchers have conducted analysis on the methods used to generate such inauthentic comments, which were far more rudimentary than is possible with today’s AI [268]. Sockpuppets continue to overwhelm US government public comment processes, and few agencies have taken action to systematically filter out or catch the malicious actors perpetrating these abuses [64, 202]. Recently, a bill aiming to curb foreign and AI-powered abuse in the public comment process for the Bureau of Land Management (the American Voices in Federal Lands Act) was introduced in the US Senate [190].
@@
By reducing the impact of sockpuppets, PHCs might also help to advance a wide range of beneficial experiments in digital governance in the public, private, and nonprofit sectors, though these are speculative and also come with risks. [^63] Scholars have argued that new democratic innovations may be especially important in the age of highly capable AI [6, 25]. For instance, AI tools might help to digitally gather and analyze public opinions from large segments of the population [98]—PHCs can help ensure that people do not use multiple personas to wield more influence and thereby can increase the perceived legitimacy of such processes. [^64] PHCs can help digital governance processes be more robust to claims that AI bots were used to influence a particular outcome. For more on the ability to cast doubt on outcomes by invoking the specter of AI manipulation, see [225]. In particular, PHCs can reduce the impact of communities that might otherwise intentionally undermine these processes. [^65] One challenge in soliciting broad input online is that particular communities might flood a poll to try to undermine the poll’s intent. For instance, a poll to name a new Mountain Dew flavor produced leading options like “Hitler did nothing wrong” [124], forcing the organizers to take down the naming process. For more, see [219]. Looking ahead, collective input on the design and deployment of AI systems is an emerging special case for soliciting representative opinions, free from sockpuppet manipulation. [^66] Some AI providers have commissioned research on how to best channel popular input in their tools, including via “Alignment Assemblies” to better align models with democratic will [282, 86, 63, 13].
@@ 3.2   Mitigate bot attacks
Malicious actors use botnets—coordinated groups of entities meant to appear as distinct users—to carry out more effective attacks online. Sometimes, overwhelming a Web service with a number of distinct-appearing users is the point of the attack, as in distributed denial-of-service (DDoS) attacks [237]. Other times, the purpose of relying upon many distinct-appearing users is to create many independent chances of success, like circumventing a service’s rules and creating new accounts for abuse even after having been caught. As discussed in Section 1, without new mitigations, malicious actors powered by AI systems will be more likely to succeed at these attacks.
@@
Personhood credentials can help reduce the impact of bot attacks by enabling service providers to suspend users whose accounts have engaged in abuse. Today, adversaries’ tactics—like changing IP addresses and spoofing different browser configurations—often successfully evade the restrictions in place, at least for a period [210]. Some existing software does guard against these attacks, but the developers of cybersecurity software have been locked in a perpetual cat-and-mouse game with attackers. And now, AI changes the dynamic, powering the attackers by reducing the manual work needed to get a new botnet running again after a previous one was caught. By using personhood credentials, service providers can restrict each individual to a limited number of accounts. This means that if someone breaks the rules and their account is suspended, they—identified through their personhood credential used at registration—cannot create new accounts without limit. [^67] When establishing restrictions on user activity, websites must balance the level of abuse they are willing to tolerate against the potential loss of legitimate user engagement that such limits might cause [167].
@@
Bot attacks have wide-ranging harms. Genuine users of peer-to-peer digital services are taxed when bots are not kept in check—they face increased friction and are targets of scams. [^68] For instance, a website may have particular processes—phone number verification, rate limits—that are primarily meant to limit abuse by repeat bad actors but that inadvertently create difficulties for benign users. Digital communities that aim to allow any individual to claim some asset (e.g., free trial memberships, credits for computational costs, products for beta testing), but only once, are often undermined by bot attacks. PHCs can serve as a valuable tool for checking who has already received the benefit, without revealing or tracking any further information. [^69] A prosaic example of such fraud is free trials for digital services, which are often serially abused [77]. This rate-limiting can also be applied to mitigate other forms of abuse that are exacerbated by rapid, AI-driven actions, such as ticket scalping, which disadvantages ordinary buyers [250, 187]. Similarly, fraudulent AI-powered requests for benefits from governments or aid-giving organizations—even when successfully defended against—can drain resources and make it more difficult to serve the targets of the aid; screening by applicants verified with PHCs may help. [^70] One risk is that screening via PHCs may reduce fraud but also make it more difficult for legitimate recipients to receive aid: Administrative overhead to apply for benefits tends to result in many people not receiving benefits that they are properly eligible for [120, 226].
@@
Looking ahead, we expect that many forms of AI-powered deceptive behavior that are especially harmful at scale can be indirectly mitigated by personhood credentials. In the vignette in Scenario 2, for example, we discuss one particular form of AI-powered deception that is harmful in isolated incidents but especially harmful at scale—impersonation. [^71] By impersonation, we mean taking on the likeness of someone in particular, as opposed to taking on a sockpuppet persona—a plausible human that could exist but does not. We dwell on this case because it nicely delineates how PHCs can and cannot directly help when it comes to AI-powered deception online: PHCs by default cannot validate a holder’s legal identity, and so a digital likeness service cannot directly check a PHC to confirm that the holder is taking on their own likeness rather than another person’s. [^72] Malicious actors may have incentives to impersonate a real person to carry out social engineering attacks or to set up a profit-making scheme (e.g., if the target of impersonation is a celebrity). However, such activities can be prohibited by digital services’ policies, and if a user is ever found to be in violation of the service’s policies, PHCs can make the policies more enforceable, by suspending the PHC-linked account and disallowing future signups for that PHC. [^73] Beyond violating the policies of a digital service, using AI for impersonation might also sometimes involve breaking a law. Courts have sometimes ruled in favor of a “right to publicity” and “right to privacy,” which using a person’s likeness without their permission may violate, depending on the circumstances [269, 101]. We expect that—anticipating these consequences—fewer attackers will attempt such unauthorized impersonations.
@@ 3.3   Allow verified delegation to AI agents
Scenarios 3 Accountable delegation to AI assistants on anonymous platforms.
@@
AI systems are increasingly capable of acting autonomously [141, 264, 95, 277]. While enabling many beneficial use cases, [^74] For a description of some economically useful properties of AI agents, see [231]. Already, some humans are deferring to AI-powered solutions for navigating dating apps on their behalf [230]. this autonomy (“agenticness”) also facilitates a new form of deception: bad actors can deploy AI systems that, instead of pretending to be a person, accurately present as AI agents but pretend to act on behalf of a user who does not exist. This strategy exploits the current lack of norms around disclosure of agentic AIs, including a lack of norms around disclosing the identities of the people controlling them (often called their “principals”). [^75] In some jurisdictions there is already regulation governing bot disclosure [143]. In 2023, legislation was introduced to codify bot-disclosure federally, though it has not become law [189, 258]. Note that these bills outline requirements around disclosure of AI usage, i.e., requiring that people disclose when a bot was employed to take an action or when a piece of content was generated by an AI model. The outlined disclosure requirements stop there—they do not include disclosures related to the identity of the person or organization enlisting the AI.
@@
Personhood credentials could offer a way to verify that AI agents are acting as delegates of real people, signaling credible supervision without revealing the principal’s legal identity. This feature could be useful in a range of settings where users wish to rely upon AI assistants; Scenario 3 outlines one such use case. Should a principal fail to address harms caused by their PHC-verified AI, they risk suspension from a service. Suspension implies that they lose their ability to verify delegates for some time period, reducing their capacity to perpetrate future harms.
@@
Note how, in this case, PHCs create a form of accountability for AI agents without demanding sensitive information from principals. [^76] A large and growing body of legal scholarship explores when and how humans should be accountable for the harms caused by AI agents [87, 54, 55, 52, 56, 156, 74, 144]. Many—though by no means all [238, 100]—of these proposals involve holding humans liable for some harms caused most directly by those agents. It is beyond the scope of this article to recommend any of these approaches over others or to explore the finer points of how such theories could work. We note, however, that these theories rely on the ability of someone harmed by an AI agent to sue the principal of that agent, which in turn depends on the principal being identifiable—which a PHC alone does not achieve. Even without directly identifying the principal, however, PHCs can still shift the benefits of AI agent usage to be more positive. This suspension mechanism may effectively signal which AI agents have trustworthy principals, even if verifying AI agents through principal-linked PHCs is voluntary. Agents that remain unverified might be perceived as having reasons for not undergoing verification. [^77] This mechanism aligns with the “unraveling result” of voluntary disclosure models in economic theory [172, 173]. In scenarios where parties hold verifiable private information—such as whether an AI agent operates under a trustworthy principal—even if revealing this information is optional, agents associated with trustworthy principals have a strong incentive to disclose it. Consequently, a lack of disclosure becomes informative: in equilibrium, agents that do not disclose are effectively signaling that they do not have trustworthy principals. Moreover, some malicious activities involving autonomous AI agents may rely on hiding the fact that multiple agents are controlled by the same individual. [^78] One example of how a person might enlist unverified agents to manipulate a market: A potential buyer on an auction site might create multiple personas to submit lowball offers, in hopes of baiting the seller into an artificially low assessment of their item’s value. In a similar example, a potential seller might enlist synthetic personas to create the appearance of higher demand for their item, inducing a legitimate buyer to pay a higher price and to act with more urgency. In each case, the ability to recognize these entities as linked to the same identity—even without knowing any further details of the identity—would help to head off these risks. For an estimation of the prevalence of these practices, known as “shill-bidding,” in online markets, see [51]. For a discussion of market designs that are robust to these practices, see [145]. PHCs can help address this issue by creating a framework that links multiple AI agents to a single principal without revealing the principal’s specific identity. By doing so, PHCs might make it more difficult for bad actors to conceal their network of AI agents, thereby reducing the potential for abuse that stems from undisclosed common principals.
@@
Sometimes, a website—or a third-party user interacting with the agent—may need to verify the AI’s specific principal, not merely that it is backed by some principal. Ultimately, a fuller framework for verifying AI agents [43, 44] and their principals will likely be necessary.